OpenStudy (anonymous):

find the solution set: |2x+1|=7

OpenStudy (anonymous):

x=3,-4

OpenStudy (anonymous):

2x+1 = 7 or 2x+1 = -7 2x+1 =7 2x = 6 x = 3 2x+1 = -7 2x = -8 x= -4

OpenStudy (anonymous):

so 3, -4

OpenStudy (anonymous):

I'm sorry, I am awful at math - can you simply explain how to do it?

OpenStudy (anonymous):

|x| = a mean x = a or x = -a now look at my solution

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

I appreciate it, it just takes me a minute to sink in & really understand. Math is a foreign language to me.